A spherical pressure vessel is constructed by bolting two hemispheres together along rigid flanges as shown in the sketch. The sphere has an internal diameter of 1.5 m and the plate thickness is 5 mm. Allowable tensile stresses are 45 MPa in the shell and 80 MPa in the bolts.

 

1      Calculate the allowable gauge pressure.                                                        

2      Determine the number of bolts required if the bolt diameter is 25 mm        

           (Note: bolt root diameter = 0.84 × nominal bolt diameter)


1
Expert's answer
2021-07-08T04:33:01-0400

Part 1

"S= \frac{PD}{2t}\\
S= \frac{45*1.5}{2*\frac{5}{1000}}\\
S= 6750 Nm^{-2}"


Part 2

"n(AS_w) = PA\\
n =\frac{ PA}{AS_w}\\
n =\frac{ 6750* 4\pi*0.75^2 }{\frac{\pi}{4} *0.025^2*80}\\
n= 121500 \space bolts"